# Client-Aided Secure Two-Party Computation of Dynamic Controllers

### v2

@@ 1 Introduction
The need for protecting privacy and security in control systems has become critical due to growing cyber-physical systems in various industries. Encrypted control, which integrates automatic control with cryptographic technology, has emerged to prevent eavesdropping attacks and securely outsource controller computations for an untrusted third party [1, 2, 3]. The initial realization of encrypted control was achieved in a single-server setting by using homomorphic encryption [4]. Homomorphic encryption is an encryption scheme that allows arithmetic operations such as addition and multiplication on encrypted data [5, 6, 7, 8, 9]. In this framework, as illustrated in Fig. 11(a), the server receives encrypted measurements from the sensor, performs controller computation using encrypted parameters without decryption, and returns encrypted control actions to the actuator. Unlike other privacy mechanisms, such as differential privacy [10], encrypted control offers reliable security built on cryptography without a tradeoff between control performance and security level [3].
With existing homomorphic encryption schemes, arithmetic operations on encrypted data are usually performed on a finite range of integers. One of the main challenges in this context is to prevent overflow in recursive state updates when encrypting dynamic controllers (see Section 3). To address the overflow issue, fully homomorphic encryption with bootstrapping was employed in [11], and the periodic reset of controller states was considered in [12]. Recent approaches to this issue involve reformulating controller representations. For instance, one approach approximated a controller using a finite impulse response filter [13]. In addition, another approach reformulated a controller based on pole placement and coordinate transformation using re-encrypted control inputs [14]. The input re-encryption was also applied in [15] and [16], which represented controller states as historical input and output data.
@@
While these methods have advanced encrypted control in a single-server setting, exploring encrypted control in a multi-party computation setting presents potential advantages. Multi-party computation using secret sharing [17] is generally more computationally efficient and secure than homomorphic encryption in exchange for an increase in communication costs. Furthermore, multi-party computation offers the ability to evaluate various functionalities. This property motivates the implementation of encrypted control in a multi-party setting because homomorphic encryption typically requires substantial computational resources for computing mathematical operations beyond addition and multiplication.
Fig. 11(b) illustrates a typical system architecture of two-party encrypted control using secret sharing. Each controller receives a share of sensor measurements, securely performs controller computation (along with communication to the other controller if needed), and returns a share of control inputs to the actuator. In this framework, an efficient and secure implementation of static controllers was suggested [18]. This method was extended to polynomial control by combining secret sharing with homomorphic encryption [19]. The polynomial control was also realized based only on secret sharing in more than a three-party setting [20]. Furthermore, in a two-party setting, secret sharing with a garbled circuit was applied to implement a neural-network-based controller [21], and model predictive control using homomorphic encryption was considered [22].
@@
  • We realize secure two-party computation of any linear time-invariant controllers without requiring state decryption or input re-encryption. This requirement of previous encrypted dynamic controls in a single-server setting [14, 15, 16] is undesirable because of increasing computation and communication efforts on a client. Compared to the previous schemes, the proposed protocol is simpler and significantly reduces the computation costs of a client.
@@ 2 Problem Setting
We adopt a client-aided computation model [23] for two-party computation. In this setting, the client who initiates queries to two parties provides not only the protocol inputs but also auxiliary inputs, although it does not interact with them during the computation. The parties use these auxiliary inputs to execute the protocol efficiently and return the outputs to the client. We also assume peer-to-peer secure channels as a network model, which can be realized by standard symmetric-key encryption. Furthermore, throughout this paper, we assume the following standard assumptions for two-party computation.
@@ 3 Dynamic Controllers Over Integers
To perform controller computation with secret sharing, the controller parameters, states, inputs, and outputs should be encoded into matrices and vectors over integers. This section revisits the well-known overflow problem [24, 14] caused by such encoding and clarifies the nature of the difficulty of encrypting dynamic controllers using homomorphic encryption.
@@
With the encoded parameters, state, and output, one might consider the following controller over integers [11, 12],
@@
To avoid this problem, one can consider another controller with truncation [14] given as
@@
This approximation suggests that truncation solves the problem of controller state overflow, making (6) suitable for encoding the controller (2). However, implementing the encoded controller using homomorphic encryption remains problematic. The difficulty lies in the fact that homomorphic encryption is supposed to be unable to compute the truncation $\lceil\cdot/2^{\ell}\rfloor$ on encrypted data directly [14]. This is a fundamental problem in the encryption of dynamic controllers using homomorphic encryption. We address this problem by constructing an efficient protocol to compute the truncation using secret sharing, further discussed in Section 5.
@@ 4 Secret Sharing and Beaver Triple
This section provides a brief review of secret sharing and its arithmetic [17, 25].
@@ 4.2 Beaver Triple
To compute valid shares of $xy\bmod q$ privately, we use a technique based on Beaver triples [25]. A Beaver triple over $\mathbb{Z}_{q}$ is a triplet $(\llbracket a\rrbracket,\llbracket b\rrbracket,\llbracket c\rrbracket)$ that consists of shares of random numbers $a,b,c\in\mathbb{Z}_{q}$ such that $c=ab\bmod q$. With the triplet, the multiplication of shares is computed as follows [25, 17]. Suppose that the parties $P_{i}$, $i\in\{1,2\}$ have $\llbracket x\rrbracket_{i}$, $\llbracket y\rrbracket_{i}$, and $(\llbracket a\rrbracket_{i},\llbracket b\rrbracket_{i},\llbracket c\rrbracket_{i})$. Each party computes $\llbracket d\rrbracket_{i}=\llbracket x\rrbracket_{i}-\llbracket a\rrbracket_{i}\bmod q$ and $\llbracket e\rrbracket_{i}=\llbracket y\rrbracket_{i}-\llbracket b\rrbracket_{i}\bmod q$ locally. The parties open $\llbracket d\rrbracket_{i}$ and $\llbracket e\rrbracket_{i}$ and compute $d=\mathsf{Reconst}(\llbracket d\rrbracket)$ and $e=\mathsf{Reconst}(\llbracket e\rrbracket)$, where $\llbracket d\rrbracket=(\llbracket d\rrbracket_{1},\llbracket d\rrbracket_{2})$ and $\llbracket e\rrbracket=(\llbracket e\rrbracket_{1},\llbracket e\rrbracket_{2})$. Then, $P_{1}$ and $P_{2}$ respectively compute $\llbracket z\rrbracket_{1}=e\llbracket a\rrbracket_{1}+d\llbracket b\rrbracket_{1}+\llbracket c\rrbracket_{1}+de\bmod q$ and $\llbracket z\rrbracket_{2}=e\llbracket a\rrbracket_{2}+d\llbracket b\rrbracket_{2}+\llbracket c\rrbracket_{2}\bmod q$ locally. The shares $\llbracket z\rrbracket=(\llbracket z\rrbracket_{1},\llbracket z\rrbracket_{2})$ satisfy $\llbracket z\rrbracket=\llbracket xy\bmod q\rrbracket$ because
@@
Under Assumption 3, the 22-out-of-22 secret sharing scheme satisfies perfect security [26]. This security means that parties can learn nothing from their own shares, even if their computational capabilities are unlimited.
@@ 5 Two-Party Computation of Dynamic Controllers
To enable this, we first introduce a truncation protocol originally proposed in [27] and slightly modified here to suit our setting, along with its corresponding security notion. We then apply the protocol to realize (6) over shares and analyze the control performance achievable by the proposed protocol.
@@ 5.1 Truncation Protocol
An important insight behind the truncation protocol in [27] is that, by definition of the modulo operation, the truncation of $\ell$ bits of $m\in\mathbb{Z}$ can be represented by
@@
The protocol takes as input not only $\llbracket m\rrbracket=(\llbracket m\rrbracket_{1},\llbracket m\rrbracket_{2})$ and $\ell$ but also shares $\llbracket r\rrbracket=(\llbracket r\rrbracket_{1},\llbracket r\rrbracket_{2})$ and $\llbracket r^{\prime}\rrbracket=(\llbracket r^{\prime}\rrbracket_{1},\llbracket r^{\prime}\rrbracket_{2})$ with random numbers $r\in\mathbb{Z}_{\langle\kappa-\ell+\lambda\rangle}$ and $r^{\prime}\in\mathbb{Z}_{\langle\ell\rangle}$. Suppose that parties $P_{i}$, $i\in\{1,2\}$ have $\llbracket m\rrbracket_{i}$, $\llbracket r\rrbracket_{i}$, and $\llbracket r^{\prime}\rrbracket_{i}$. Each party computes $\llbracket m_{r}\rrbracket_{i}=\llbracket m\rrbracket_{i}+2^{\ell}\llbracket r\rrbracket_{i}+\llbracket r^{\prime}\rrbracket_{i}+2^{\ell-1}\bmod q$ locally. The party $P_{2}$ transmits $\llbracket m_{r}\rrbracket_{2}$ to $P_{1}$, and $P_{1}$ reconstructs $m_{r}=\mathsf{Reconst}(\llbracket m_{r}\rrbracket)$, where $\llbracket m_{r}\rrbracket=(\llbracket m_{r}\rrbracket_{1},\llbracket m_{r}\rrbracket_{2})$. Then, $P_{1}$ and $P_{2}$ locally compute $\llbracket m^{\prime}\rrbracket_{1}=\operatorname{inv}(2^{\ell},q)(\llbracket m\rrbracket_{1}+\llbracket r^{\prime}\rrbracket_{1}-(m_{r}-2^{\ell-1}\bmod 2^{\ell}))\bmod q$ and $\llbracket m^{\prime}\rrbracket_{2}=\operatorname{inv}(2^{\ell},q)(\llbracket m\rrbracket_{2}+\llbracket r^{\prime}\rrbracket_{2})\bmod q$, respectively. Consequently, $\llbracket m^{\prime}\rrbracket=(\llbracket m^{\prime}\rrbracket_{1},\llbracket m^{\prime}\rrbracket_{2})$ becomes shares of $\lceil m/2^{\ell}\rfloor+w$ for all $m\in\mathbb{Z}_{\langle\kappa\rangle}$, where $w\in\{-1,0,1\}$.
@@
The lemma shows that the $\mathsf{Trunc}$ protocol truncates the least $\ell$ bits of $m\in\mathbb{Z}_{\langle\kappa\rangle}$ with a $1$-bit error $w$. This error occurs only when $m+r^{\prime}\notin\mathbb{Z}_{\langle\ell\rangle}$ and thus can be removed by an additional protocol to detect $m+r^{\prime}\notin\mathbb{Z}_{\langle\ell\rangle}$ over shares [27]. The following section tolerates the error to avoid increasing the computational and communication costs and applies the truncation protocol to approximate and implement the encoded controller (6) to run it over shares.
@@
The security of $\mathsf{Trunc}$ is identical to that of the truncation protocol in [27]. Through protocol execution, $P_{2}$ does not receive any messages from $P_{1}$ and thus learns no information. The only information leaked into $P_{1}$ is $m_{r}=m+2^{\ell}r+r^{\prime}+2^{\ell-1}\bmod q$. Under the conditions of bit lengths of $m$, $r$, and $r^{\prime}$, it holds that $m_{r}=m+r^{\prime\prime}\in\mathbb{Z}_{q}$, where $r^{\prime\prime}=2^{\ell}r+r^{\prime}+2^{\ell-1}$ is uniformly distributed over $\mathbb{Z}_{\langle\kappa+\lambda\rangle}$. Intuitively, $m_{r}$ is a masked message of $m$ by $r^{\prime\prime}$, and thus $P_{1}$ cannot learn any information from $m_{r}$. Indeed, the statistical distance between $m+r^{\prime\prime}$ and $r^{\prime\prime}$, $|\frac{1}{2}\sum_{x\in\mathbb{Z}_{q}}\absolutevalue*{\Pr[m+r^{\prime\prime}=x]-\Pr[r^{\prime\prime}=x]}$, is less than $2^{-\lambda}$ and so is negligible in $\lambda$. Therefore, the $\mathsf{Trunc}$ protocol achieves statistical security, which is a relaxation of the perfect security. However, statistical security is still valid for unconditional adversaries. This implies that it is stronger than computational security achieved by homomorphic encryption [5, 6, 7, 8, 9]. Note that computational security covers only conditional adversaries whose computational capability is limited to polynomial time [28].

### v3

@@ 1 Introduction
The need for protecting privacy and security in control systems has become critical due to growing cyber-physical systems in various industries. Encrypted control, which integrates automatic control with cryptographic technology, has emerged to prevent eavesdropping attacks and securely outsource controller computations for an untrusted third party [Darup2021-qq, Kim2022-ck, Schluter2023-ia]. The initial realization of encrypted control was achieved in a single-server setting by using homomorphic encryption [Kogiso2015-go]. Homomorphic encryption is an encryption scheme that allows arithmetic operations such as addition and multiplication on encrypted data [Regev2009-ys, Brakerski2014-ny, Fan2012-jt, Gentry2013-eq, Cheon2017-pd]. In this framework, as illustrated in Fig. 11(a), the server receives encrypted measurements from the sensor, performs controller computation using encrypted parameters without decryption, and returns encrypted control actions to the actuator. Unlike other privacy mechanisms, such as differential privacy [Han2018-pm], encrypted control offers reliable security built on cryptography without a tradeoff between control performance and security level [Schluter2023-ia].
With existing homomorphic encryption schemes, arithmetic operations on encrypted data are usually performed on a finite range of integers. One of the main challenges in this context is to prevent overflow in recursive state updates when encrypting dynamic controllers (see Section 3). To address the overflow issue, fully homomorphic encryption with bootstrapping was employed in [Kim2016-lc], and the periodic reset of controller states was considered in [Murguia2020-ov]. Recent approaches to this issue involve reformulating controller representations. For instance, one approach approximated a controller using a finite impulse response filter [Schluter2021-ek]. In addition, another approach reformulated a controller based on pole placement and coordinate transformation using re-encrypted control inputs [Kim2023-pk]. The input re-encryption was also applied in [Teranishi2024-ha] and [Lee2025-jo], which represented controller states as historical input and output data.
@@
While these methods have advanced encrypted control in a single-server setting, exploring encrypted control in a multi-party computation setting presents potential advantages. Multi-party computation using secret sharing [Cramer2015-sj] is generally more computationally efficient and secure than homomorphic encryption in exchange for an increase in communication costs. Furthermore, multi-party computation offers the ability to evaluate various functionalities. This property motivates the implementation of encrypted control in a multi-party setting because homomorphic encryption typically requires substantial computational resources for computing mathematical operations beyond addition and multiplication.
Fig. 11(b) illustrates a typical system architecture of two-party encrypted control using secret sharing. Each controller receives a share of sensor measurements, securely performs controller computation (along with communication to the other controller if needed), and returns a share of control inputs to the actuator. In this framework, an efficient and secure implementation of static controllers was suggested [Darup2019-ou]. This method was extended to polynomial control by combining secret sharing with homomorphic encryption [Darup2020-cm]. The polynomial control was also realized based only on secret sharing in more than a three-party setting [Schlor2021-am]. Furthermore, in a two-party setting, secret sharing with a garbled circuit was applied to implement a neural-network-based controller [Tjell2021-ar], and model predictive control using homomorphic encryption was considered [Alexandru2018-jv].
@@
  • We realize secure two-party computation of any linear time-invariant controllers without requiring state decryption or input re-encryption. This requirement of previous encrypted dynamic controls in a single-server setting [Kim2023-pk, Teranishi2024-ha, Lee2025-jo] is undesirable because of increasing computation and communication efforts on a client. Compared to the previous schemes, the proposed protocol is simpler and significantly reduces the computation costs of a client.
@@ 2 Problem Setting
We adopt a client-aided computation model [Ohata2020-py] for two-party computation. In this setting, the client who initiates queries to two parties provides not only the protocol inputs but also auxiliary inputs, although it does not interact with them during the computation. The parties use these auxiliary inputs to execute the protocol efficiently and return the outputs to the client. We also assume peer-to-peer secure channels as a network model, which can be realized by standard symmetric-key encryption. Furthermore, throughout this paper, we assume the following standard assumptions for two-party computation.
@@ 3 Dynamic Controllers Over Integers
To perform controller computation with secret sharing, the controller parameters, states, inputs, and outputs should be encoded into matrices and vectors over integers. This section revisits the well-known overflow problem [Cheon2018-vr, Kim2023-pk] caused by such encoding and clarifies the nature of the difficulty of encrypting dynamic controllers using homomorphic encryption.
@@
With the encoded parameters, state, and output, one might consider the following controller over integers [Kim2016-lc, Murguia2020-ov],
@@
To avoid this problem, one can consider another controller with truncation [Kim2023-pk] given as
@@
This approximation suggests that truncation solves the problem of controller state overflow, making (6) suitable for encoding the controller (2). However, implementing the encoded controller using homomorphic encryption remains problematic. The difficulty lies in the fact that homomorphic encryption is supposed to be unable to compute the truncation $\lceil\cdot/2^{\ell}\rfloor$ on encrypted data directly [Kim2023-pk]. This is a fundamental problem in the encryption of dynamic controllers using homomorphic encryption. We address this problem by constructing an efficient protocol to compute the truncation using secret sharing, further discussed in Section 5.
@@ 4 Secret Sharing and Beaver Triple
This section provides a brief review of secret sharing and its arithmetic [Cramer2015-sj, Beaver2007-kr].
@@ 4.2 Beaver Triple
To compute valid shares of $xy\bmod q$ privately, we use a technique based on Beaver triples [Beaver2007-kr]. A Beaver triple over $\mathbb{Z}_{q}$ is a triplet $(\llbracket a\rrbracket,\llbracket b\rrbracket,\llbracket c\rrbracket)$ that consists of shares of random numbers $a,b,c\in\mathbb{Z}_{q}$ such that $c=ab\bmod q$. With the triplet, the multiplication of shares is computed as follows [Beaver2007-kr, Cramer2015-sj]. Suppose that the parties $P_{i}$, $i\in\{1,2\}$ have $\llbracket x\rrbracket_{i}$, $\llbracket y\rrbracket_{i}$, and $(\llbracket a\rrbracket_{i},\llbracket b\rrbracket_{i},\llbracket c\rrbracket_{i})$. Each party computes $\llbracket d\rrbracket_{i}=\llbracket x\rrbracket_{i}-\llbracket a\rrbracket_{i}\bmod q$ and $\llbracket e\rrbracket_{i}=\llbracket y\rrbracket_{i}-\llbracket b\rrbracket_{i}\bmod q$ locally. The parties open $\llbracket d\rrbracket_{i}$ and $\llbracket e\rrbracket_{i}$ and compute $d=\mathsf{Reconst}(\llbracket d\rrbracket)$ and $e=\mathsf{Reconst}(\llbracket e\rrbracket)$, where $\llbracket d\rrbracket=(\llbracket d\rrbracket_{1},\llbracket d\rrbracket_{2})$ and $\llbracket e\rrbracket=(\llbracket e\rrbracket_{1},\llbracket e\rrbracket_{2})$. Then, $P_{1}$ and $P_{2}$ respectively compute $\llbracket z\rrbracket_{1}=e\llbracket a\rrbracket_{1}+d\llbracket b\rrbracket_{1}+\llbracket c\rrbracket_{1}+de\bmod q$ and $\llbracket z\rrbracket_{2}=e\llbracket a\rrbracket_{2}+d\llbracket b\rrbracket_{2}+\llbracket c\rrbracket_{2}\bmod q$ locally. The shares $\llbracket z\rrbracket=(\llbracket z\rrbracket_{1},\llbracket z\rrbracket_{2})$ satisfy $\llbracket z\rrbracket=\llbracket xy\bmod q\rrbracket$ because
@@
Under Assumption 3, the 22-out-of-22 secret sharing scheme satisfies perfect security [Shannon1949-mc]. This security means that parties can learn nothing from their own shares, even if their computational capabilities are unlimited.
@@ 5 Two-Party Computation of Dynamic Controllers
To enable this, we first introduce a truncation protocol originally proposed in [Escudero2020-ih] and slightly modified here to suit our setting, along with its corresponding security notion. We then apply the protocol to realize (6) over shares and analyze the control performance achievable by the proposed protocol.
@@ 5.1 Truncation Protocol
An important insight behind the truncation protocol in [Escudero2020-ih] is that, by definition of the modulo operation, the truncation of $\ell$ bits of $m\in\mathbb{Z}$ can be represented by
@@
The protocol takes as input not only $\llbracket m\rrbracket=(\llbracket m\rrbracket_{1},\llbracket m\rrbracket_{2})$ and $\ell$ but also shares $\llbracket r\rrbracket=(\llbracket r\rrbracket_{1},\llbracket r\rrbracket_{2})$ and $\llbracket r^{\prime}\rrbracket=(\llbracket r^{\prime}\rrbracket_{1},\llbracket r^{\prime}\rrbracket_{2})$ with random numbers $r\in\mathbb{Z}_{\langle\kappa-\ell+\lambda\rangle}$ and $r^{\prime}\in\mathbb{Z}_{\langle\ell\rangle}$. Suppose that parties $P_{i}$, $i\in\{1,2\}$ have $\llbracket m\rrbracket_{i}$, $\llbracket r\rrbracket_{i}$, and $\llbracket r^{\prime}\rrbracket_{i}$. The parties compute $\llbracket m_{r}\rrbracket=\llbracket m\rrbracket+2^{\ell}\llbracket r\rrbracket+\llbracket r^{\prime}\rrbracket+2^{\ell-1}$ locally. The party $P_{2}$ transmits $\llbracket m_{r}\rrbracket_{2}$ to $P_{1}$, and $P_{1}$ reconstructs $m_{r}=\mathsf{Reconst}(\llbracket m_{r}\rrbracket)$, where $\llbracket m_{r}\rrbracket=(\llbracket m_{r}\rrbracket_{1},\llbracket m_{r}\rrbracket_{2})$. Then, $P_{1}$ and $P_{2}$ locally compute $\llbracket m^{\prime}\rrbracket_{1}=\operatorname{inv}(2^{\ell},q)(\llbracket m\rrbracket_{1}+\llbracket r^{\prime}\rrbracket_{1}-(m_{r}-2^{\ell-1}\bmod 2^{\ell}))\bmod q$ and $\llbracket m^{\prime}\rrbracket_{2}=\operatorname{inv}(2^{\ell},q)(\llbracket m\rrbracket_{2}+\llbracket r^{\prime}\rrbracket_{2})\bmod q$, respectively. Consequently, $\llbracket m^{\prime}\rrbracket=(\llbracket m^{\prime}\rrbracket_{1},\llbracket m^{\prime}\rrbracket_{2})$ becomes shares of $\lceil m/2^{\ell}\rfloor+w$ for all $m\in\mathbb{Z}_{\langle\kappa\rangle}$, where $w\in\{-1,0,1\}$.
@@
The lemma shows that the $\mathsf{Trunc}$ protocol truncates the least $\ell$ bits of $m\in\mathbb{Z}_{\langle\kappa\rangle}$ with a $1$-bit error $w$. This error occurs only when $m+r^{\prime}\notin\mathbb{Z}_{\langle\ell\rangle}$ and thus can be removed by an additional protocol to detect $m+r^{\prime}\notin\mathbb{Z}_{\langle\ell\rangle}$ over shares [Escudero2020-ih]. The following section tolerates the error to avoid increasing the computational and communication costs and applies the truncation protocol to approximate and implement the encoded controller (6) to run it over shares.
@@
The security of $\mathsf{Trunc}$ is identical to that of the truncation protocol in [Escudero2020-ih]. Through protocol execution, $P_{2}$ does not receive any messages from $P_{1}$ and thus learns no information. The only information leaked into $P_{1}$ is $m_{r}=m+2^{\ell}r+r^{\prime}+2^{\ell-1}\bmod q$. Under the conditions of bit lengths of $m$, $r$, and $r^{\prime}$, it holds that $m_{r}=m+r^{\prime\prime}\in\mathbb{Z}_{q}$, where $r^{\prime\prime}=2^{\ell}r+r^{\prime}+2^{\ell-1}$ is uniformly distributed over $\mathbb{Z}_{\langle\kappa+\lambda\rangle}$. Intuitively, $m_{r}$ is a masked message of $m$ by $r^{\prime\prime}$, and thus $P_{1}$ cannot learn any information from $m_{r}$. Indeed, the statistical distance between $m+r^{\prime\prime}$ and $r^{\prime\prime}$, $|\frac{1}{2}\sum_{x\in\mathbb{Z}_{q}}\absolutevalue*{\Pr[m+r^{\prime\prime}=x]-\Pr[r^{\prime\prime}=x]}$, is less than $2^{-\lambda}$ and so is negligible in $\lambda$. Therefore, the $\mathsf{Trunc}$ protocol achieves statistical security, which is a relaxation of the perfect security. However, statistical security is still valid for unconditional adversaries. This implies that it is stronger than computational security achieved by homomorphic encryption [Regev2009-ys, Brakerski2014-ny, Fan2012-jt, Gentry2013-eq, Cheon2017-pd]. Note that computational security covers only conditional adversaries whose computational capability is limited to polynomial time [Katz2014-kb].